2-3 January 2009

Observers: J.Hill (Tucson) and A. Rakich (standing by)
Software Support: C. Biddick (Tucson)
Telescope Operator: J.Rousseau (mtn) plus J. Hill (Tucson)

Summary

Had some minor telescope problems at startup. The most important one is that we had to jumper out the LUCIFER lockout switch for the rotator.

The new DSP slewing algorithm has been tested during they day, and now for half a dozen slews at night.

We got a star image for a few minutes. So at least GCS v2.0 has now been tested briefly on sky.

Then we had to close for increasing humidity and fog.

Details

Dave A. has given us a new and improved version of MCSPU with the new command pre-processor (no jumps) and with slewing managed directly by the DSPs rather than by the MCSPU.

We also have updated versions of PSF, and a version of GCS that hasn't been previously tested on sky.

Open

17:50 Open all 4 vent doors.

18:15 Start Mirror Ventilation. Right side MV gave a PB0402 fail to run alarm. Jon R. reset the breaker. Mirror Temps are +2 degC.

18:10 Handover from Dave A's MCS engineering.

18:15 We had a problem that both PMC subsystems were hung. Chris restarted them and now they seem OK.

18:20 Start MCSPU (after Dave logged off).

18:20 Open Shutter doors. Close side Vent Doors. Close rear Vent Doors to midway. T=-2.0 degC D=-6.0 degC Wind 10 m/sec. Sky is almost photometric with a few clouds far in the southwest.

18:40 We get a mysterious instance of "Lock Out" on the telescope AZ and EL axes. Jon R. is the only one there, so we know that nobody turned the key. It seems to have been a spurious display on the MCSPU engineering interface. elready and azready work as normal.

18:55 LFBG Rotator also appears to be locked. A simple rcready lfbg doesn't cure the problem, so may it really has some issue.

18:57 elslewtohold 90 works but prints "this is a default error message".

19:05 Jon resets the rotator MCR button, but that doesn't help. Both the left chain and the left rotator are reporting hardware faults, and the LOCK is reported LOCKED, and the MCR is reported OPEN.

19:10 Cycling power on the left rotator drive has not helped.

19:20 Call Dave A. to see if he can understand the rotator problem. It appears that the new lock out circuit that runs into the LUCIFER electronics box is open. We'll have it going once Jon installs a bypass jumper.

19:25 Enable Actuator 430 on DX mirror.

19:48 pointto BS9116 gs=0 ACQUIRE (still too low) but we never got the movie window from GCSGUI.

19:54 pointto M5_0182 ACQUIRE

19:55 Stopped and restarted GCSL. But GCS crashed a few minutes later. Restarted.

19:56 pointto M5_0182 ACQUIRE

A getStatus command seems to have very slow communications between GCS and oacserver. (10's of seconds)

20:01 Rebooted agw-control to get a fresh version of oacserver. Now getStatus runs at normal speed.

20:05 pointto M5_0182 ACQUIRE failed because couldn't move probe

20:06 stopAGW / startAGW

20:07 pointto M5_0182 ACQUIRE still couldn't move probe (not homed)

20:08 home the AGW motors

20:09 pointto M5_0182 ACQUIRE (warning about hitting M1 Y limit since I forgot to clear active optics, although the image wasn't so bad)

20:11 Clear Active Optics. Adjust IE=+2 CA=+6 (defaults were IE=-15 CA=-11)

20:17 pointto BS9108 gs=0 ACTIVE - Now doing active optics with PSF

Close

20:27 Active Optics has dialed the image down to 0.7 arcsec, and the humidity is rising rapidly with fog appearing outside. slew telescope to horizon Close Shutter Doors; Close Vent Doors; Stop Mirror Ventilation.

Had problems with the Ball Detents on Latch Pin 1007.

Open

22:16 Humidity is back below 80%, so we reopen. Again an alarm on Latch Pin 1007, and on PB0402.

22:23 pointto BS9108 gs=0 ACTIVE - couldn't find star (there was no star on the acquisition image 00002)

22:26 pointto M5_0182 ACQUIRE - star was centered on ACQUIRE movie image to within 2 arcsec. Why didn't we find previous star?

22:29 pointto BS9107 gs=0 ACTIVE - succeeded (star was in the extreme upper right corner of the acquisition image 00003)

Note that both this preset, and the one from 20:17 waited for 2 good guide iterations before returning from the preset command (as seen in the GCS and PCS syslogs).

Close

22:30 Humidity is back to 95.6% (SMT) so we close again.

Guiding Tests

Field Measurements

Sky Maps: M5 Stars

Sky Maps: Person's Standards

Star Charts

Person's Star =
Start time =

Guide Star x y PA Doug's Angle wfscimage000*
- 0 612.5 0 9 1-3
1 0       4-6
- 0 612.5 90 -81 7-9
1 0       10-12

End time =

-- DougMiller - 01 Jan 2009

Topic revision: r10 - 10 Mar 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback